OpenStudy (anonymous):

On a certain day, the company took 88 people on whale watching trips. There were 44 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (anonymous):

(44 + x)35 + (88 – (44 + x)) ∙ 46 = C (88 - x)35 + (88 – 44) ∙ 46 = C (44 – x)35 + (88 – 44) ∙ 46 = C (88 + x)35 + (88 – (44 + x)) ∙ 46 = C

OpenStudy (anonymous):

those r the choices right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the answers b

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

yea ive done this question b4

OpenStudy (anonymous):

The ANSWER IS C

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

no fare for children below 3 yrs (makes sense right) fare for children between 3 and 12 is 35 fare for people above 12 is 46 hence the final answer would be \[x*0 + (44-x)*35 + (88-44)*46 = C(total money collected)\]